# KnowsLM: A framework for evaluation of small language models for knowledge augmentation and humanised conversations

### v1

@@ Abstract
In the evolving landscape of conversational AI, generating concise, context-aware, and human-like dialogue using small and medium-sized language models (LLMs) remains a complex challenge. KnowsLM introduces a unique framework to help evaluate enhancement of Small and medium Language Models with domain-specific knowledge and stylistically adaptive conversational abilities through different techniques viz. fine-tuning and retrieval-augmented generation (RAG)A medium-sized language model was fine-tuned on diverse synthetic datasets covering region-specific food suggestions, poetic response styles, cultural contexts such as the MahaKumbh, and recent events like ISRO’s space docking mission. Our study investigates the influence of LoRA rank, dataset scale, and prompt prefix design on both knowledge retention and stylistic alignment. While fine-tuning improves fluency and enables stylistic customization, its ability to integrate unseen knowledge is constrained—particularly with smaller datasets. Conversely, RAG-augmented models, equipped to incorporate external documents at inference, demonstrated superior factual accuracy on out-of-distribution prompts, though they lacked the stylistic consistency achieved by fine-tuning. Evaluations by LLM-based judges across knowledge accuracy, conversational quality, and conciseness suggest that fine-tuning is best suited for tone adaptation, whereas RAG excels at real-time knowledge augmentation. Together, these insights inform practical strategies for deploying small and medium LLMs in scalable, human-centric dialogue systems.
The KnowsLM framework showcases the complementary benefits of fine-tuning and RAG—where fine-tuning enhances coherence and stylistic consistency within familiar contexts, and RAG effectively injects external or novel information. This study offers practical guidance for developing efficient, cost-effective methods to augment small and medium LLMs for generating more human-like, context-aware conversations.
@@ 4 Conclusion
KnowsLM presents an approach to enrich small and medium-sized language models with external knowledge and concise, human-like conversational responses. It demonstrates the effective use of retrieval-augmented generation (RAG) for dynamic knowledge integration, alongside fine-tuning on synthetically generated dialogue pairs derived from source datasets to infuse stylistic and tonal authenticity.Our experiments underscore the nuanced trade-offs involved in enhancing LLaMA 3.3 70B through fine-tuning and retrieval-augmented strategies. While local and stylistic adaptations proved effective in aligning model outputs with cultural and tonal expectations, the benefits were contingent on both data quality and scale. RAG-based augmentation offered clear advantages in factual accuracy, whereas fine-tuning better preserved conversational fluency. However, fine-tuning on small datasets yielded limited gains, with significant improvements only emerging at larger scales—albeit with increased computational cost. These findings highlight the critical balance between dataset size, model performance, and resource efficiency in building robust domain-specific language models.

### v2

@@ Abstract
In the evolving landscape of conversational AI, generating concise, context-aware, and human-like dialogue using small and medium-sized language models (LLMs) remains a complex challenge. KnowSLM introduces a unique framework to help evaluate enhancement of Small and Medium Language Models with domain-specific knowledge and stylistically adaptive conversational abilities through different techniques viz. Fine-tuning and retrieval-augmented generation (RAG). The KnowSLM framework showcases the complementary benefits of fine-tuning and RAG, where fine-tuning enhances coherence and stylistic consistency within familiar contexts, and RAG effectively injects external or novel information. KnowSLM framework has been applied on a Medium language model, which has been fine-tuned on various synthetic datasets comprising of region-specific food suggestions, poetic response styles, cultural contexts such as MahaKumbh, and recent events such as ISRO’s space docking mission. Our study investigates the influence of LoRA rank, dataset scale, and prompt prefix design on both knowledge retention and stylistic alignment. Although fine-tuning improves fluency and enables stylistic customization, its ability to integrate unseen knowledge is restricted, particularly with smaller datasets. In contrast, RAG-augmented models, equipped to incorporate external documents at inference, demonstrated superior factual accuracy on out-of-distribution prompts, though they lacked the stylistic consistency achieved by fine-tuning. Evaluations by LLM-based judges on knowledge accuracy, conversational quality, and conciseness suggest that fine-tuning is best suited for tone adaptation, whereas RAG excels at real-time knowledge enhancement. Together, these insights inform practical strategies for deploying small and medium LLMs in scalable, human-centric dialogue systems.
@@ 4 Conclusion
KnowSLM presents an approach to enrich small and medium-sized language models with external knowledge and concise, human-like conversational responses. It demonstrates the effective use of retrieval-augmented generation (RAG) for dynamic knowledge integration, alongside fine-tuning on synthetically generated dialogue pairs derived from source datasets to infuse stylistic and tonal authenticity. Our experiments underscore the nuanced trade-offs involved in enhancing LLaMA 3.3 70B through fine-tuning and retrieval-augmented strategies. We observed that local and stylistic adaptations proved effective in aligning model outputs with cultural and tonal expectations, the benefits were contingent on both data quality and scale. RAG-based augmentation offered clear advantages in factual accuracy, on the other hand fine-tuning better preserved conversational fluency. However, fine-tuning on small datasets yielded limited gains, with significant improvements only emerging at larger scales—albeit with increased computational cost. These findings highlight the critical balance between dataset size, model performance, and resource efficiency in building robust domain-specific language models.